OpenStudy (toxicsugar22):

For the following reaction, 6.99 grams of hydrogen gas are mixed with excess nitrogen gas. The reaction yields 27.7 grams of ammonia. N2 (g) + 3 H2 (g) = 2 NH3 (g) (1) What is the theoretical yield of ammonia ? grams (2) What is the percent yield for this reaction ? %

OpenStudy (toxicsugar22):

What is the Maximum amount of Agl, in grams, that can be produced in the reaction Amount Nal reacting = 0.348 Theoetical Agl formed = 0.348

OpenStudy (toxicsugar22):

a student conduct an expirement by reacting 52.1 g Nal with an excess of AgNO3. The reaction yields 60.468 Agl. What is the percent yield. AgNO3 (aq)+ Nal(aq)= AgI(s)+NaNO3(aq)

OpenStudy (toxicsugar22):

Can you help me aswer this

OpenStudy (anonymous):

First moles of hydrogen =6.99/2 3 H2= 2 NH3 6.99/2= X MOLES NH3, which is 2.33 moles moles of NH3=mass/MM of NH3 2.33=mass/17 theoritical mass 0f NH3 =2.33*17=39.61 grams A)39.61,theoritical value as you can see from the equation above so called theory B)percent yield=(actual /theoritical)*100 now actual =27.7 G theoritical =39.61G from a baove therefore P%=(27.7/39.61)*100=69.93183%

OpenStudy (anonymous):

i am sorry this is the answer 69.9318%

OpenStudy (anonymous):

please i dont think a percent yield can actually be more than a hundred

OpenStudy (anonymous):

moles of NaI = 52.1/23+127 which is 325.625 E-3 MASS OF AgI = 325.625E-3 * (196+127) which is 105.176875 grams which is the theoritical value the actual value is 60.468 therefore % yield= (60.468/105.176875 ) *100 which is 57.4917%

OpenStudy (anonymous):

pleaase when i dont calculate well ,forgive me it is because i get tired sometimes from anatomy i am reading

OpenStudy (toxicsugar22):

OHH YOUR IN ANATOMY mE TO

OpenStudy (toxicsugar22):

what school do you go to